OpenStudy (anonymous):

Boyle's Law states that when a sample of gas is compressed at a constant temperature, the pressure P and the volume V satisfy the equation PV=C, where C is a constant. Suppose that at a certain instant the volume is 600cm3, the pressure is 150kPa, and the pressure is increasing at a rate of 20kPa/min. At what rate is the volume decreasing at this instant?

OpenStudy (anonymous):

please help i got -80cm^3/min

OpenStudy (anonymous):

I plugged Im my numbers for the problem into this Solve for the constant C first. 200(600)=C C=120000 If C is a constant, you can rewrite the equation an V= C/P. Take the derivative and find that dV= -C/(P^2) dP by using the quotient rule for calculus. Now plug in your initial values dV= -(120000)/(200)^2 *10 dV= - 30 cm^3/min.

OpenStudy (anonymous):

ok so how would i use that to make my problem work?